OpenStudy (anonymous):

A projectile has maximum range 200m. What is the max height attained by it?

OpenStudy (anonymous):

Do you have the formulas for range and height?

OpenStudy (anonymous):

is the answer 50?

OpenStudy (anonymous):

Are you simply checking it?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Would you care to show your working?

OpenStudy (anonymous):

ok max range=v^2/g=200m therefore v=sqrt(2000) subs in eq for max height (vsin$)^2/2g for $=45 degree g =10m/s^2 i get max height 50m is it correct?

OpenStudy (anonymous):

Yep, correct. Therefore, to conclude, for an ideal projectile motion, \[R\max = 4 H\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

and thats Hmax.

OpenStudy (anonymous):

noted